OpenStudy (anonymous):

choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. A: y = -6x - 15 B:y = -6x - 20 C:y = -6x + 15 D:y = -6x + 20

OpenStudy (notamathgenius):

So we knows its going to be negative of course XD what else do we know

OpenStudy (anonymous):

Lol idk:P

OpenStudy (anonymous):

SO a or b?

OpenStudy (notamathgenius):

I hate this kind of math but working together here it will help

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

WOuld is be A?

OpenStudy (notamathgenius):

If you think A yea why not XD math sucks the other one i did with you was easy this one not so easy

OpenStudy (anonymous):

HAha

OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x - 2y = 8. x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

Can you do this one?

OpenStudy (notamathgenius):

Ok so you really need to do is plug in the numbers for x and y and if you get 8 YOU WIN

OpenStudy (anonymous):

>.< That's hard :(

OpenStudy (anonymous):

Is the answer A?

OpenStudy (notamathgenius):

its not hard its confussing :D ill draw it

OpenStudy (anonymous):

D?

OpenStudy (notamathgenius):

use a calculator to solve it :D

OpenStudy (anonymous):

They all equal 6 >.<

OpenStudy (notamathgenius):

really -_- wow stupid math hmmm

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (notamathgenius):

well after figuring them all are 6 no -_- i would say guess but thats only a dire need choice